can someone please help me?
Amy can buy 3 sandwiches and 4 iced teas for $22.25. She can buy 1 sandwich and 3 iced teas for $9.50. Find the cost for 1 sandwich and 1 iced tea.

Answer by geetha_rama(94)

Let x be the cost of sandwich and y be the cost of tea
3x + 4y = 22.25 -> equation 1
x + 3y = 9.5 -> equation 2
Multiple equation 2 by 3
=> 3x + 9y = 28.5
3x + 4y = 22.25 (equation 1)

Subtract the above two equations,
5y = 6.25
y = 1.25
x = 9.5 - 3y ( from equation 2)
=9.5 - 3.75
= 5.75
Cost of 1 Sandwich = $5.75
Cost of 1 iced tea = $1.25